An issue was encountered with the load balancers for the outbound SMTP mail service which resulted in a small proportion of connections being intermittently refused, resulting in mail queuing on your server(s). As a result users may have noticed a delay in their email being delivered to recipients as it… Continue Reading Outbound SMTP load balancing
Earlier this evening we encountered an outage impacting vlinweb4. At this time an engineer is investigating the root cause, following reinstatement of service. At this time it is believed the outage occurred due to be a memory leak. Investigations are focusing on determine the specific source.
Routine monitoring has highlighted intermittent access issues for customers with Windows based hosting account. We encountered an issue during the past hour where there was intermittent access to email hosting for customers with Windows based hosting. We have resolved the stability issue and normal access is now restored. We will… Continue Reading Email stability (resolved)
A number of services encountered a short period of downtime this evening due to a SAN controller freezing. We have restarted the controller and service has now been restored. We will now begin a detailed investigation to understand the issue fully and prevent the issue from occurring again.
Our monitoring has identified an issue with upstream connectivity and our data centre provider is currently investigating. Intermittent access to services is likely to be experienced at present, an update will be posted shortly as more information becomes available.
We experienced an unplanned service outage between approximately10am and 11am this morning which affected our ‘shared services’ . 90% of services were restored by 11am with the remaining 10% of services restored subsequently. The precise nature of the issue is being investigated and further details will be placed here within approximately 24… Continue Reading Shared serice outage 06/09/2012
We experienced stability issues with web server WEB2 over the past 3 hours which resulted in a period where web pages were struggling to be served. We have isolated the issue and restored normal service as of 02:00. A period of monitoring will continue to ensure no further stability issues.
Our monitoring has highlighted a packet loss issue which is impacting our external connectivity. Customers may see this as slow or intermittent access to services. We are currently in contact with our upstream network provider.
We experienced a network issue this evening on one of our cluster nodes this evening at approximately 23:00 which resulted in some websites and email being unavailable for approximately 10 minutes. Further investigations will be conducted to isolate the cause of the crash and the reason why automated recovery did not occur.
Our monitoring has identified an issue exchanging traffic with some third party networks. We are currently investigationg this issue.